OpenStudy (anonymous):

The solution to the radical function √x+3=2 is x = 8. True False

OpenStudy (anonymous):

False. √x+3=2 √x =2-3 x = -1^2 x = 1

OpenStudy (anonymous):

You my dear, are amazing!

OpenStudy (anonymous):

Just basic questions you should get hang of it now :D.

OpenStudy (anonymous):

you have no idea how cruddy I am at math :P